OpenStudy (anonymous):

What is C in x^2 + x = 12

OpenStudy (anonymous):

-12. put it in standard form and set it equal to 0

OpenStudy (anonymous):

yeah, what's C supposed to stand for?

OpenStudy (yttrium):

take note of the general form of a quadratic equation. \[ax^2 + bx + C = 0\] in your case it is \[x ^2+x=\] so you have to make it \[x^2+x-12 = 0\] btw, c refers to constant term

OpenStudy (yttrium):

correction on the second equation \[x^2+x = 12\]

OpenStudy (anonymous):

x^2 + x -12=0 standard from is ax^2 + bx + c = 0 a b and c are constants (the opposite of a variable) that are actual numbers. x^2 is like writing 1x^2 and that 1 is what a would be.

OpenStudy (anonymous):

well what if I do -x^2 - x + 12 = 0? now C = 12.

OpenStudy (yttrium):

well in reality, we make the leading coefficient as positive. that why we have to make it x^2+x-12=0